To solve quadratic equations-
The sum of three positive, consecutive numbers is 149. find the numbers.
letx=the first number

Answer by checkley77(12844)

Let x, x+1 & x+2 be the three numbers.
x+(x+1)+(x+2)=149
3x+3=149
3x=149-3
3x=146
x=146/3
x=48.6667 ans for the first number.
48.6667+1=49.6667 ans for the second number.
48.6667+2=50.6667 ans for the third number.
Proof:
48.6667+49.6667+50.6667=149
149~149